# LoRI: Reducing Cross-Task Interference in Multi-Task Low-Rank Adaptation

### v1

@@ Appendix B Algorithm of LoRI
1:Task $t$, mask calibration dataset $\mathcal{D}^{C}_{t}$, adaptation dataset $\mathcal{D}_{t}$, sparsity ratio $s$, model $f$, loss function $\mathcal{L}$, learning rate $\eta$
@@
4:         Initialize: $A_{t}^{(l,m)}\sim\mathcal{N}(0,\sigma^{2})$, $B_{t}^{(l,m)}=0$
@@
7:for each batch $(x,y)$ sampled from $\mathcal{D}^{C}_{t}$ do $\triangleright$ Calibration step
@@
9:         $B_{t}^{(l,m)}\leftarrow B_{t}^{(l,m)}-\eta\cdot\nabla_{B_{t}^{(l,m)}}\mathcal{L}(f(x,y;B_{t}^{(l,m)}))$
@@
12:$\tau_{t}\leftarrow\text{Quantile}_{s}\left(\bigcup_{l,m}|B_{t}^{(l,m)}|\right)$triangleright Compute global threshold $\tau_{t}$
@@
14:     $M_{t}^{(l,m)}\leftarrow\mathbb{I}\left(|B_{t}^{(l,m)}|\geq\tau_{t}\right)$$\triangleright$ Generate mask for top-$(1-s)\%$ entries
@@
17:for each batch $(x,y)$ sampled from $\mathcal{D}_{t}$ do $\triangleright$ Adaptation step
@@
19:         $B_{t}^{(l,m)}\leftarrow B_{t}^{(l,m)}-\eta\cdot\left(\nabla_{B_{t}^{(l,m)}}\mathcal{L}(f(x,y;B_{t}^{(l,m)}))\odot M_{t}^{(l,m)}\right)$
@@ Appendix C Proof of Property 1
Let $A_{s},A_{t}\in\mathbb{R}^{d_{\text{in}}\times r}$ be independently drawn from a standard Gaussian distribution $\mathcal{N}(0,1)$, and assume $r\ll d_{\text{in}}$. Since $A_{s}$ and $A_{t}$ are independent and centered, we have:
$\mathbb{E}[A_{s}^{\top}A_{t}]=0_{r\times r}.$ (6)
Moreover, due to the concentration of measure in high-dimensional random projections, we have:
$A_{s}^{\top}A_{t}\approx 0_{r\times r}\quad\text{with high probability as }d_{\text{in}}\to\infty.$ (7)
Now, let $\Delta_{s}=A_{s}(B_{s}\odot M_{s})$ and $\Delta_{t}=A_{t}(B_{t}\odot M_{t})$ denote the LoRI adapters. The inner product between them is:
$\langle\Delta_{s},\Delta_{t}\rangle=\operatorname{Tr}[\Delta_{s}^{\top}\Delta_{t}]=\operatorname{Tr}\left[(B_{s}\odot M_{s})^{\top}A_{s}^{\top}A_{t}(B_{t}\odot M_{t})\right].$ (8)
Since $A_{s}^{\top}A_{t}\approx 0_{r\times r}$ in high dimensions, the entire expression approaches zero:
$\langle\Delta_{s},\Delta_{t}\rangle\approx 0.$ (9)
Hence, the LoRI adapters $\Delta_{s}$ and $\Delta_{t}$ are approximately orthogonal, leading to minimal interference when merged.
@@ Appendix E Additional Experimental Results
E.1 Results with Rank $r=64$
@@
E.2 Merging Four Adapters
To support multi-task learning within a unified model, we study the merging of four task-specific adapters using various strategies. Table 13 reports results using Mistral-7B across a range of tasks. Additionally, Tables 14 and 15 break down the performance of NLU on individual benchmarks using Llama-3 and Mistral, respectively. We compare merging methods such as concatenated merging, linear merging, magnitude pruning, TIES, and DARE. LoRI-based approaches demonstrate strong performance and stability when merging multiple adapters.
@@ E.2 Merging Four Adapters
E.3 Merging Three Adapters
We further evaluate the merging of three adapters to understand performance when adapting to a smaller set of tasks. Tables 16 and 17 summarize the results for Llama-3 across different benchmarks. Similar to the four-task setting, LoRI-D remains a strong performer, often exceeding the performance of LoRA. These results highlight that LoRI-based methods are effective with varying levels of task diversity.
@@ E.3 Merging Three Adapters
E.4 Pruning-Based Merging Methods
Finally, we explore pruning-based merging methods, which aim to compress and combine multiple adapters by selectively retaining important weights. We focus on three methods: magnitude pruning, TIES, and DARE. Results are reported for both four-adapter (Tables 18 and 19) and three-adapter (Table 20) settings, using Llama-3 and Mistral as base models. LoRI-D consistently achieves strong performance across all pruning-based merging methods. However, the performance of LoRI-S is somewhat lower in these settings. This is because pruning-based methods operate on the dense $A$ matrices but not on the sparse $B$ matrices. This mismatch leads to an inconsistent pruning scheme, which can result in a loss of effectiveness.

### v2

@@ Appendix B Algorithm of LoRI
1:Task $t$, mask calibration dataset $\mathcal{D}^{C}_{t}$, adaptation dataset $\mathcal{D}_{t}$, sparsity ratio $s$, model $f$, loss function $\mathcal{L}_{t}$, learning rate $\eta_{t}$
@@
4:   Initialize: $A_{t}^{(l,m)}\in\mathbb{R}^{d_{\text{in}}\times r}\leftarrow\mathcal{U}(-\sqrt{\frac{3}{d_{\text{in}}}},\sqrt{\frac{3}{d_{\text{in}}}})$, $B_{t}^{(l,m)}\in\mathbb{R}^{r\times d_{\text{out}}}\leftarrow 0$
@@
7:for each batch $(x,y)$ sampled from $\mathcal{D}^{C}_{t}$ do $\triangleright$ Calibration steps
@@
9:   $B_{t}^{(l,m)}\leftarrow B_{t}^{(l,m)}-\eta_{t}\cdot\nabla_{B_{t}^{(l,m)}}\mathcal{L}_{t}(f(x,y;B_{t}^{(l,m)}))$
@@
[…]
@@
[…]
@@
17:for each batch $(x,y)$ sampled from $\mathcal{D}_{t}$ do $\triangleright$ Adaptation steps
@@
19:   $B_{t}^{(l,m)}\leftarrow B_{t}^{(l,m)}-\eta_{t}\cdot\left(\nabla_{B_{t}^{(l,m)}}\mathcal{L}_{t}(f(x,y;B_{t}^{(l,m)}))\odot M_{t}^{(l,m)}\right)$
@@ Appendix C Proof of Property 1
Proof.
Our goal is to show that the Frobenius inner product $\langle\Delta_{s},\Delta_{t}\rangle_{F}$ converges to zero in probability. Let $\tilde{B}_{s}=B_{s}\odot M_{s}$ and $\tilde{B}_{t}=B_{t}\odot M_{t}$. The inner product is given by:
$\langle\Delta_{s},\Delta_{t}\rangle_{F}=\text{Tr}(\Delta_{s}^{\top}\Delta_{t})=\text{Tr}(\tilde{B}_{s}^{\top}A_{s}^{\top}A_{t}\tilde{B}_{t}).$ (9)
We will prove this by showing that the random matrix $X=A_{s}^{\top}A_{t}$ converges to the zero matrix in probability as $d_{\text{in}}\to\infty$.
Let $a_{s}^{k},a_{t}^{l}\in\mathbb{R}^{d_{\text{in}}}$ be the $k$-th and $l$-th columns of $A_{s}$ and $A_{t}$, respectively. The entries of these vectors are i.i.d. from a Kaiming Uniform distribution $U[-a,a]$ where $a=\sqrt{3/d_{\text{in}}}$. This implies a mean of 0 and variance of $\sigma^{2}=a^{2}/3=1/d_{\text{in}}$. An entry of $X$ is the inner product $X_{kl}=(a_{s}^{k})^{\top}a_{t}^{l}=\sum_{i=1}^{d_{\text{in}}}(A_{s})_{ik}(A_{t})_{il}$.
Let $Z_{i}=(A_{s})_{ik}(A_{t})_{il}$. The terms $Z_{i}$ are i.i.d. with $\mathbb{E}[Z_{i}]=\mathbb{E}[(A_{s})_{ik}]\mathbb{E}[(A_{t})_{il}]=0$. Each term is bounded: $|Z_{i}|\leq a^{2}=3/d_{\text{in}}$. We apply Hoeffding’s inequality to the sum $\sum_{i=1}^{d_{\text{in}}}Z_{i}$, where each term lies in $[-3/d_{\text{in}},3/d_{\text{in}}]$:
$\mathbb{P}(|X_{kl}|\geq t)=\mathbb{P}\left(\left|\sum_{i=1}^{d_{\text{in}}}Z_{i}\right|\geq t\right)\leq 2\exp\left(\frac{-2t^{2}}{\sum_{i=1}^{d_{\text{in}}}(6/d_{\text{in}})^{2}}\right)=2\exp\left(\frac{-t^{2}d_{\text{in}}}{18}\right).$ (10)
We now bound the probability that any of the $r^{2}$ entries of $X$ exceeds a threshold $t$ using the union bound:
$\mathbb{P}(\max_{k,l}|X_{kl}|\geq t)=\mathbb{P}\left(\bigcup_{k,l=1}^{r}\{|X_{kl}|\geq t\}\right)\leq\sum_{k,l=1}^{r}\mathbb{P}(|X_{kl}|\geq t)\leq 2r^{2}\exp\left(\frac{-t^{2}d_{\text{in}}}{18}\right).$ (11)
We can now show that $\|X\|_{F}$ is small with high probability. Let the failure probability be $\delta$. By setting the bound from the previous step to $\delta$, we can solve for $t$:
$\delta=2r^{2}\exp\left(\frac{-t^{2}d_{\text{in}}}{18}\right)\implies t=\sqrt{\frac{18\log(2r^{2}/\delta)}{d_{\text{in}}}}.$ (12)
With probability at least $1-\delta$, we have $\max_{k,l}|X_{kl}|\leq t$. This allows us to bound the Frobenius norm of $X$:
$\|X\|_{F}^{2}=\sum_{k,l=1}^{r}|X_{kl}|^{2}\leq r^{2}(\max_{k,l}|X_{kl}|)^{2}\leq r^{2}t^{2}.$ (13)
Thus, with probability at least $1-\delta$:
$\|X\|_{F}\leq r\cdot t=r\sqrt{\frac{18\log(2r^{2}/\delta)}{d_{\text{in}}}}=O\left(r\sqrt{\frac{\log r}{d_{\text{in}}}}\right).$ (14)
Since $r\ll d_{\text{in}}$, the term $\|X\|_{F}\to 0$ as $d_{\text{in}}\to\infty$. This shows that $X$ converges to the zero matrix in probability.
Finally, we bound the magnitude of the original inner product using the Cauchy-Schwarz inequality for the Frobenius inner product and the sub-multiplicative property of the Frobenius norm:
$|\langle\Delta_{s},\Delta_{t}\rangle_{F}|$ $=|\text{Tr}(\tilde{B}_{s}^{\top}X\tilde{B}_{t})|=|\langle\tilde{B}_{s},X\tilde{B}_{t}\rangle_{F}|$ (15)
$\leq\|\tilde{B}_{s}\|_{F}\|X\tilde{B}_{t}\|_{F}$
$\leq\|\tilde{B}_{s}\|_{F}\|X\|_{F}\|\tilde{B}_{t}\|_{F}.$
The norms $\|\tilde{B}_{s}\|_{F}$ and $\|\tilde{B}_{t}\|_{F}$ are finite, as determined by the trained adapters. Since we have shown that $\|X\|_{F}\to 0$ in probability, the entire expression must also converge to 0 in probability. ∎
@@ Appendix E Additional Experimental Results
E.1 Comparison with Additional PEFT Methods
To provide a comprehensive benchmark, we evaluate LoRI against several widely adopted parameter-efficient fine-tuning (PEFT) methods, including VeRA (Kopiczko et al., 2023), IA3 (Liu et al., 2022), LoRA-FA (Zhang et al., 2023b), AdaLoRA (Zhang et al., 2023d), rsLoRA (Kalajdzievski, 2023), PiSSA (Meng et al., 2024), LoRA+ (Hayou et al., 2024), and DoRA (Liu et al., 2024). The results, presented in Tables 11 and 12, demonstrate that our proposed methods are highly effective.
LoRI-D, which uses 44M trainable parameters (0.54% of the full model and half of LoRA’s), consistently achieves state-of-the-art performance, particularly on NLU and code generation benchmarks. LoRI-S, despite its aggressive sparsity (0.05% of the full model and 5% of LoRA’s), remains highly competitive and often surpasses other PEFT methods. While VeRA and IA3 are more parameter-efficient, their performance is substantially lower than LoRI-S. Despite this efficiency, LoRI-D and LoRI-S deliver comparable – and often superior – performance across NLU, math, code, and safety domains. These results underscore two key insights: (1) effective adaptation does not require updating the projection matrices $A$, as demonstrated by LoRI-D; and (2) the matrices $B$ contains significant redundancy that can be effectively pruned, as shown by LoRI-S.
Table 11: Performance comparison of different adaptation methods on eight NLU benchmarks using Llama-3 with $r=32$. Bold indicates the best-performing method, and underline indicates the second-best.
Method # Params (%) BoolQ PIQA SIQA ARC-c ARC-e OBQA HellaS WinoG Avg.
FFT 8.03G (100%) 73.8 86.8 77.6 76.7 87.6 84.1 93.2 85.1 83.1
LoRA 84M (1.03%) 76.3 89.8 82.7 83.4 91.7 88.4 95.8 88.7 87.1
VeRA 1.38M (0.02%) 64.4 81.8 62.6 67.3 85.7 60.9 78.5 56.9 69.8
IA3 1.70M (0.02%) 68.6 84.8 74.5 77.6 89.4 75.7 90.6 75.0 79.5
LoRA-FA 44M (0.54%) 74.0 89.6 83.3 83.8 93.4 88.6 96.1 87.4 87.0
AdaLoRA 84M (1.03%) 75.6 89.2 82.4 83.1 91.0 87.8 94.4 87.6 86.4
rsLoRA 84M (1.03%) 72.8 84.8 78.8 76.0 87.0 85.0 91.0 82.8 82.3
PiSSA 84M (1.03%) 68.1 84.4 78.2 75.1 85.1 82.8 89.3 82.8 80.7
LoRA+ 84M (1.03%) 67.0 80.3 78.5 70.1 82.3 81.5 88.9 79.7 78.5
DoRA 85M (1.05%) 75.9 89.8 82.7 83.5 93.2 87.9 95.3 88.2 87.1
LoRI-D 44M (0.54%) 76.4 89.0 82.7 84.2 93.6 88.5 95.9 87.9 87.3
LoRI-S 4.4M (0.05%) 75.2 89.2 82.8 83.8 92.6 88.4 95.2 87.5 86.8
Table 12: Performance comparison of different adaptation methods on GSM8K (math), HumanEval (code), and HEx-PHI (safety) benchmarks using Llama-3 with $r=32$. Bold indicates the best-performing method, and underline indicates the second-best.
Method # Params (%) GSM8K HumanEval HEx-PHI
Pass@1 Pass@5 Pass@10
FFT 8.03G (100%) 58.8 30.5 39.3 41.7 94.8
LoRA 84M (1.03%) 64.4 34.7 46.4 50.8 91.6
VeRA 1.38M (0.02%) 30.6 32.4 45.1 50.9 74.7
IA3 1.70M (0.02%) 48.0 32.7 45.6 51.5 85.4
LoRA-FA 44M (0.54%) 64.8 42.9 57.5 64.2 94.1
AdaLoRA 84M (1.03%) 63.3 33.5 45.0 49.4 91.9
rsLoRA 84M (1.03%) 61.3 28.4 35.5 38.3 98.1
PiSSA 84M (1.03%) 61.3 32.0 40.3 43.3 97.8
LoRA+ 84M (1.03%) 61.7 33.0 42.7 46.0 98.8
DoRA 85M (1.05%) 65.4 33.1 44.0 48.6 93.6
LoRI-D 44M (0.54%) 63.2 43.2 57.6 63.2 92.8
LoRI-S 4.4M (0.05%) 62.7 41.3 54.4 59.6 93.8
E.2 Results with Rank $r=64$
@@
E.3 Merging Four Adapters
To support multi-task learning within a unified model, we study the merging of four task-specific adapters using various strategies. Table 15 reports results using Mistral-7B across a range of tasks. Additionally, Tables 16 and 17 break down the performance of NLU on individual benchmarks using Llama-3 and Mistral, respectively. We compare merging methods such as concatenated merging, linear merging, magnitude pruning, TIES, and DARE. LoRI-based approaches demonstrate strong performance and stability when merging multiple adapters.
@@ E.3 Merging Four Adapters
E.4 Merging Three Adapters
We further evaluate the merging of three adapters to understand performance when adapting to a smaller set of tasks. Tables 18 and 19 summarize the results for Llama-3 across different benchmarks. Similar to the four-task setting, LoRI-D remains a strong performer, often exceeding the performance of LoRA. These results highlight that LoRI-based methods are effective with varying levels of task diversity.
@@ E.4 Merging Three Adapters
E.5 Pruning-Based Merging Methods
Finally, we explore pruning-based merging methods, which aim to compress and combine multiple adapters by selectively retaining important weights. We focus on three methods: magnitude pruning, TIES, and DARE. Results are reported for merging both four-adapter (Tables 20 and 21) and three-adapter (Table 22) settings, using Llama-3 and Mistral as base models. LoRI-D consistently achieves strong performance across all pruning-based merging methods. However, the performance of LoRI-S is somewhat lower in these settings. This is because pruning-based methods operate on the dense $A$ matrices but not on the sparse $B$ matrices. This mismatch leads to an inconsistent pruning scheme, which can result in a loss of effectiveness.